OpenStudy (anonymous):

Solve this with Quad. formula please? x^2 – 7x – 6 = 0

OpenStudy (anonymous):

\[\frac{-b \pm \sqrt{49-4(1)(-6)}}{ 2 }\]

OpenStudy (anonymous):

can u simplify from there?

OpenStudy (anonymous):

srry plug in -b for positive 7

OpenStudy (anonymous):

This is the simplified version right?: 7+(73)/2=40. and 7-(73)/2=-33